OpenStudy (anonymous):

A jar contains 5 blue marbles, 8 red marbles, 4 white marbles, and 3 purple marbles. Suppose you pick a marble at random without looking. Find P (blue or purple). Write your answer as a fraction in simplest form

OpenStudy (anonymous):

that would be the (number of blue + number of purple) / total number

OpenStudy (anonymous):

\[\frac{5+3}{5+8+4+3}=\frac{2}{5} \]